# Integer Lattice Gas with a sampling collision operator for the fluctuating Navier-Stokes Equation

### v1

@@ IV.3 Deviation from the Analytical Solution
Looking at how the lattice Boltzmann and the lattice gas algorithms behave under extreme initial conditions for the SOD shock tube system allows us to compare how both algorithms deviate from the analytical solution. As the analytical solution found by Negro is a solution to the Euler equations, it is necessarily inviscid, and thus assumes that the viscosity of the system is zero Negro et al. (2019). Therefore, the analytical solution for the SOD shock tube should differ from the data from both algorithms to some degree for any finite value of $\omega<2$, as can be seen in Fig. 8.
@@
Figure 10: The $\bar{N}$ (a), $\bar{J}$ (b), and $\bar{\pi}$ (c) moment graphs of the isothermal SOD shock tube at 75 iterations for $\omega=0.1$ comparing the lattice gas results averaged over 25,000 seeds and the lattice Boltzmann results. All systems were initialized and run identically to those in Fig. 8 aside from the differing $\omega$-value. (a) and (b) include identical analytical solutions to those in Fig 8. (c) includes the equilibrium $\bar{\pi}^{0}$ given by Eq. (70).
@@
Fig. 10 compares the lattice Boltzmann and lattice gas results for a system with a much higher viscosity, being run at $\omega=0.1$. Specifically it compares the $\bar{N}$ (Fig. 10 (a)), $\bar{J}$ (Fig. 10 (b)), and post-collision $\bar{\pi}$ (Fig. 10 (c)) moments to the applicable analytical solutions in Fig. 10 (a) and (b), and the equilibrium average $\bar{\pi}$-value in Fig. 10 (c).
@@
Though the agreement with the analytical solutions for $\bar{N}$ and $\bar{J}$ is poor due to the large difference in viscosity between the simulations and the analytical solution, the lattice gas and lattice Boltzmann results show excellent agreement with each other. Similarly, the fact that $\omega=0.1$ means that the system will not fully relax to equilibrium in one timestep, meaning that the post-collision $\bar{\pi}$-values from the lattice Boltzmann and lattice gas algorithms do not recover the equilibrium $\bar{\pi}$-value, however the two simulations agree well with each other. This is important to note, as it implies that even the highly non-equilibrium state that the SOD shock tube represents still has a Boltzmann limit that is very well approximated by the BGK entropic lattice Boltzmann algorithm, something that Fig. 2 shows is not guaranteed.
@@
To further test the deviation from the analytical solution of the lattice Boltzmann and lattice gas results, we can consider an extreme initial density disparity between $N_{l}$ and $N_{r}$. This leads to a deviation from the analytical solution because as the initial density disparity between the two regions increases, the analytical solution requires $u_{\mathrm{shock}}$-values that approach or exceed $1$. Because $1$ is the maximum possible value of $u$ a lattice Boltzmann or lattice gas system can have, simulations in this range typically show significant deviations. To show this, we consider the extreme density ratio of $20:1$ for $N_{l}$ and $N_{r}$. For this particular density ratio, the $u_{\mathrm{shock}}$-value of the analytical solution is approximately $0.9$ on the central density plateau, as opposed to the analytical solution for the density ratio from Fig. 8, which has a $u_{\mathrm{shock}}$-value on the central density plateau of only around $0.2$.
@@
Figure 11: The $\bar{N}$ (a) and $\bar{J}$ (b) moment graphs of the isothermal SOD shock tube at 75 iterations for $\omega=1$ comparing the lattice gas results averaged over 25,000 seeds and the lattice Boltzmann results. The systems were initialized with $\bar{N}$ values of $N_{l}=1,000$ and $N_{r}=50$, but were otherwise initialized and run identically to those in Fig. 8. Analytical solutions calculated identically to those in Fig. 8, aside from the different $N_{l}$ and $N_{r}$ values used, are included for $\bar{N}$ and $\bar{J}$.
@@
Looking at the velocity on the plateau of the wave ($x_{c}\leq x<x_{s}$) which is where it is maximal, there is a clear difference between the simulation results and the analytical solution. The $\bar{J}$-value on the plateau corresponds to a $u_{\mathrm{shock}}$-value of approximately $0.78$ for the lattice Boltzmann and lattice gas simulations. This velocity disparity also explains why the values of $x_{c}$ and $x_{s}$ differ between the simulated data and the analytical solutions, as the $\theta$-values for the high velocity state will not be $1/3$, leading to artifacts in the Navier-Stokes equation as was discussed previously Blommel and Wagner (2018).
@@ IV.4 Non-equilibrium Fluctuations in the Isothermal SOD Shock Tube
Utilizing the same set of simulations from Fig. 8, we examined the fluctuation of the moments in the lattice gas, comparing their average non-equilibrium fluctuations to those of global equilibrium. As the equilibrium fluctuations in $N$- and $\pi$-values are given by Eq. (31), we normalize the average non-equilibrium fluctuations for $N$ and $\pi$ with $\bar{N}$ and $\bar{\pi}$ respectively.
The equilibrium fluctuations in the $J$-moment do not have the local $\bar{J}$-value as their average, as $J$ is represented by the difference between two Poisson distributed quantities (those being $n_{1}$ and $n_{-1}$). Therefore, its global equilibrium fluctuations are Skellam distributed Skellam (1946). A Skellam distribution has a variance corresponding to the sum of the mean values of the two Poisson distributions that make up the Skellum distribution Jose (1956). The average value for our $J$-fluctuations, therefore, becomes $\bar{\pi}$.
Figure 12: Normalized nonequilibrium fluctuations of the $N$, $J$, and $\pi$ moments after 75 iterations. Initialized identically to the system in Fig. 8, the system was run for 75 iterations at $\omega=1$ and averaged over 25,000 seeds. Each moment’s variance was then calculated and normalized by dividing by its mean. For the $N$-moment this mean is the local $\bar{N}$-value (Fig. 8(a)) and for the $J$ and $\pi$ moments this mean is the $\bar{\pi}$-value from Eq. (70).
The fluctuation results can be seen in Fig. 12. In the bulk of the shock wave and the regions initialized in equilibrium, the non-equilibrium fluctuations match the equilibrium fluctuations, deviating slightly due to limited averaging. However, in the highly non-equilibrium parts of the shock wave the non-equilibrium fluctuations clearly deviate from the equilibrium fluctuations, with a slight decrease of the fluctuations in the rarefaction fan and a drastic increase around the shock front, mirroring the effects seen in the pre-collision $\pi$-values in Fig. 9. While this correlation is very evocative we don’t yet understand why it occurs.
@@
Figure 13: Normalized nonequilibrium $N$-moment fluctuations for Poisson and delta distributed initial fluctuations after 75 iterations. The Poisson distributed system was initialized identically to the system in Fig. 8, however the flat system was initialized utilizing constant $n_{i}$ values corresponding to the values of $f^{0}_{i}$ as in Eq. (80). The systems were run, and the variances were calculated and normalized identically to those in Fig. 12.
@@ IV.5 Galilean Invariance of the SOD Shock Tube
As stated previously, the non-Galilean invariance of the collision operator for Boolean lattice gases was a major factor in the widespread adoption of the BGK lattice Boltzmann model Frisch et al. (1987); Qian et al. (1992). Due to this, it is important that we show to what extent our lattice gas algorithm maintains Galilean invariance if we seek the sampling lattice gas algorithm to be competitive with the lattice Boltzmann algorithm.
In order to numerically test the Galilean invariance of our system, we initialized a set of SOD shock tube systems similarly to those in Fig. 8 for both the lattice Boltzmann and lattice gas algorithms with a lattice size of $L=1,000$ and a net velocity of zero, and a second set of identical systems with a non-zero constant velocity offset $u_{g}$. After running all simulations (including averaging 2,500 seeds for each lattice gas system) for 400 iterations at $\omega=1$, we adjusted the coordinate systems for the initial $u_{g}$-values to take into account the additional distance the simulation would have traveled on average due to the velocity by using,
@@
which takes into account our system’s periodic boundary conditions. $[a]\bmod[b]$ is the modulo operator, returning the positive remainder when $b$ divides $a$. We then calculated the mean squared error in $\bar{N}$ and $u$,
@@
A comparison of the $\epsilon$-values for $\bar{N}$ and $u$ for both the lattice gas and lattice Boltzmann algorithms are shown in Fig. 14 (a) and 14 (b) for various values of $u_{g}$ ranging from $-0.2$ to $0.2$. The mean squared error for the lattice gas algorithm is on an order of magnitude comparable to that of the lattice Boltzmann algorithm for both $\bar{N}$ and $u$, meaning that the lattice gas and lattice Boltzmann algorithms are approximately equivalent in terms of Galilean invariance. There is a slight deviation observed for higher values of $u_{g}$, which implies that the lattice gas algorithm is slightly more Galilean invariant than the lattice Boltzmann algorithm, however we are as of yet unsure why this deviation occurs.
@@ V Computational Efficiency
As the main drawback of Blommel and Wagner’s method was its lack of computational efficiency, determining the efficiency of our own method is necessary for comparison. This first requires some more technical discussion of the methods we used to make our sampling method more efficient. Because the local equilibrium ensemble out of which we wish to sample is not a well known distribution, using a library such as the GNU Scientific Library (GSL) in C is non-trivial Galassi et al. (2018). Though packages such as the GSL have the ability to sample from arbitrary distributions, the full distribution must be known in advance, which is not the case for our system. This is due to the fact that the $N$- and $J$-values that define the local $\pi$ equilibrium ensemble are not known until after the streaming step of the prior timestep. The ideal solution to this issue would be a library of full distributions pre-calculated for every combination of $N$- and $J$-values, however this method requires either a large amount of memory (scaling with $O(N)$) or to read the distribution from a file for each collision.
Thus, we decided to utilize a dynamic lookup table. This method is similar to what was done in our prior paper on the diffusive sampling lattice gas, only with the added need to manage the $J$-value as well as the $N$-value Seekins and Wagner (2022). We first initialized the lookup table as a two-dimensional triangular array with $12N-$ 1 entries in each sub-array, with a likely maximum amount of particles that would be at one lattice site given by three times the average system density. Should a state occur beyond this maximum density, the table would be rebuilt with a higher maximum density in the algorithm. As the system was running, the approximate center point of the cumulative equilibrium distribution ($\sum_{\pi=J}P^{0}(\pi;N,J)\approx 0.5$) of the local equilibrium ensemble was saved in the lookup table. This then allowed us to then navigate to any part of the local equilibrium ensemble by utilizing the recursive formula from Eq. (27) and its inverse. Because this dynamic lookup table adds entries on the fly, over the course of longer runs generating new table entries will become rare. A copy of the code for this algorithm is available on Github at Seekins (2025).
Even with this lookup table, the limiting factor of efficiency for the lattice gas is the generation of sampled random numbers. In order to investigate the maximum possible improvement in efficiency due to this sampling, we wished to determine the speed of the algorithm without the sampling as a factor. To do this, we pre-generated a full set of random numbers for a simulation, and then fed those into the system in place of the dynamic lookup table.
@@
Figure 15: Computational efficiency of the lattice Boltzmann algorithm compared to various lattice gas algorithms for average system densities from $N=10$ to $N=1,000$. (a) compares the lattice Boltzmann algorithm to Blommel and Wagner’s algorithm, our sampling lattice gas algorithm, and the maximal potential efficiency of our sampling lattice gas algorithm for $\omega=0.5$. (b) compares the lattice Boltzmann algorithm to our sampling lattice gas algorithm and its maximal efficiency for $\omega=1$, leaving out Blommel and Wagner’s algorithm since it has no true way to reach $\omega=1$. See text for further details.
We ran our efficiency tests for four separate algorithms: a lattice Boltzmann algorithm using the entropic BGK collision operator, Blommel and Wagner’s algorithm, our sampling lattice gas algorithm, and the maximal lattice gas algorithm. The results of these comparisons are shown in Fig. 15 in millions of lattice updates per second (MLUPS). Because Blommel and Wagner’s algorithm is based around individual two-particle collisions, the particle collision fraction can be estimated by taking Seekins and Wagner (2022),

### v2

@@ IV.3 Deviation from the Analytical Solution
Looking at how the lattice Boltzmann and the lattice gas algorithms behave under extreme initial conditions for the SOD shock tube system allows us to compare how both algorithms deviate from the analytical solution. As the analytical solution found by Negro is a solution to the Euler equations, it is necessarily inviscid, and thus assumes that the viscosity of the system is zero [39]. Therefore, the analytical solution for the SOD shock tube should differ from the data from both algorithms to some degree for any finite value of $\omega<2$, as can be seen in Fig. 8.
@@
Figure 10: The $\bar{N}$ (a), $\bar{J}$ (b), and $\bar{\pi}$ (c) moment graphs of the isothermal SOD shock tube at 75 iterations for $\omega=\omega_{LB}=0.1$ comparing the lattice gas results averaged over 25,000 random seeds and the lattice Boltzmann results. All systems were initialized and run identically to those in Fig. 8 aside from the $\omega$ and $\omega_{LB}$ values. (a) and (b) include identical analytical solutions to those in Fig 8. (b) also includes an inset for the Mach number calculated using Eq. (82). (c) includes the equilibrium $\bar{\pi}^{0}$ given by Eq. (72).
@@
Fig. 10 compares the lattice Boltzmann and lattice gas results for a system with a much higher viscosity, being run at $\omega=\omega_{LB}=0.1$. Specifically it compares the $\bar{N}$ (Fig. 10 (a)), $\bar{J}$ (Fig. 10 (b)), and post-collision $\bar{\pi}$ (Fig. 10 (c)) moments to the applicable analytical solutions in Fig. 10 (a) and (b), and the equilibrium average $\bar{\pi}$ value in Fig. 10 (c).
@@
Though the agreement with the analytical solutions for $\bar{N}$ and $\bar{J}$ is poor due to the large difference in viscosity between the simulations and the analytical solution, the lattice gas and lattice Boltzmann results show excellent agreement with each other. Similarly, the fact that $\omega=\omega_{LB}=0.1$ means that the system will not fully relax to equilibrium in one timestep, meaning that the post-collision $\bar{\pi}$ values from the lattice Boltzmann and lattice gas algorithms do not recover the equilibrium $\bar{\pi}$ value, however the two simulations agree well with each other. This is important to note, as it implies that even the highly non-equilibrium state that the SOD shock tube represents still has a Boltzmann limit that is very well approximated by the BGK entropic lattice Boltzmann algorithm, something that Fig. 2 shows is not guaranteed.
@@
To further test the deviation from the analytical solution of the lattice Boltzmann and lattice gas results, we can consider an extreme initial density disparity between $N_{l}$ and $N_{r}$. This leads to a deviation from the analytical solution because as the initial density disparity between the two regions increases, the analytical solution requires $u_{\mathrm{shock}}$ values that approach or exceed $1$. Because $1$ is the maximum possible value of $u$ a lattice Boltzmann or lattice gas system can have, simulations in this range typically show significant deviations. To show this, we consider the extreme density ratio of $20:1$ for $N_{l}$ and $N_{r}$. For this particular density ratio, the $u_{\mathrm{shock}}$ value of the analytical solution is approximately $0.9$ on the central density plateau, as opposed to the analytical solution for the density ratio from Fig. 8, which has a $u_{\mathrm{shock}}$ value on the central density plateau of only around $0.2$.
@@
Figure 11: The $\bar{N}$ (a) and $\bar{J}$ (b) moment graphs of the isothermal SOD shock tube at 75 iterations for $\omega=\omega_{LB}=1$ comparing the lattice gas results averaged over 25,000 random seeds and the lattice Boltzmann results. The systems were initialized with $\bar{N}$ values of $N_{l}=1,000$ and $N_{r}=50$, but were otherwise initialized and run identically to those in Fig. 8. Analytical solutions calculated identically to those in Fig. 8, aside from the different $N_{l}$ and $N_{r}$ values used, are included for $\bar{N}$ and $\bar{J}$. (b) also includes an inset for the Mach number calculated using Eq. (82)
@@
Looking at the velocity on the plateau of the wave ($x_{c}\leq x<x_{s}$) which is where it is maximal, there is a clear difference between the simulation results and the analytical solution. The $\bar{J}$ value on the plateau corresponds to a $u_{\mathrm{shock}}$ value of approximately $0.78$ for the lattice Boltzmann and lattice gas simulations. This velocity disparity also explains why the values of $x_{c}$ and $x_{s}$ differ between the simulated data and the analytical solutions, as the $\theta$ values for the high velocity state will not be $1/3$, leading to artifacts in the Navier-Stokes equation as was discussed previously [19].
@@ IV.4 Non-equilibrium Fluctuations in the Isothermal SOD Shock Tube
Utilizing the same set of simulations from Fig. 8, we examined the fluctuation of the moments in the lattice gas, comparing their average non-equilibrium fluctuations to those of global equilibrium. As the equilibrium fluctuations in $N$ and $\pi$ values are given by Eq. (30), we normalize the average non-equilibrium fluctuations for $N$ and $\pi$ with $\bar{N}$ and $\bar{\pi}$ respectively.
The equilibrium fluctuations in the $J$ moment do not have the local $\bar{J}$ value as their average, as $J$ is represented by the difference between two Poisson distributed quantities (those being $n_{1}$ and $n_{-1}$). Therefore, its global equilibrium fluctuations are Skellam distributed [40]. A Skellam distribution has a variance corresponding to the sum of the mean values of the two Poisson distributions that make up the Skellum distribution [41]. The average value for our $J$ fluctuations, therefore, becomes $\bar{\pi}$.
Figure 12: Normalized nonequilibrium fluctuations of the $N$, $J$, and $\pi$ moments after 75 iterations. Initialized identically to the system in Fig. 8, the system was run for 75 iterations at $\omega=1$ and averaged over 25,000 random seeds. Each moment’s variance was then calculated and normalized by dividing by its mean. For the $N$ moment this mean is the local $\bar{N}$ value (Fig. 8(a)) and for the $J$ and $\pi$ moments this mean is the $\bar{\pi}$ value from Eq. (72).
The fluctuation results can be seen in Fig. 12. In the bulk of the shock wave and the regions initialized in equilibrium, the non-equilibrium fluctuations match the equilibrium fluctuations, deviating slightly due to limited averaging. However, in the highly non-equilibrium parts of the shock wave the non-equilibrium fluctuations clearly deviate from the equilibrium fluctuations, with a slight decrease of the fluctuations in the rarefaction fan and a drastic increase around the shock front, mirroring the effects seen in the pre-collision $\pi$ values in Fig. 9. While this correlation is very evocative we don’t yet understand why it occurs.
@@
Figure 13: Normalized nonequilibrium NN moment fluctuations for Poisson and delta distributed initial fluctuations after 75 iterations. The Poisson distributed system was initialized identically to the system in Fig. 8, however the flat system was initialized utilizing constant $n_{i}$ values corresponding to the values of $f^{0}_{i}$ as in Eq. (83). The systems were run, and the variances were calculated and normalized identically to those in Fig. 12.
@@ IV.5 Galilean Invariance of the SOD Shock Tube
As stated previously, the non-Galilean invariance of the collision operator for Boolean lattice gases was a major factor in the widespread adoption of the BGK lattice Boltzmann model [1, 5]. Due to this, it is important that we show to what extent our lattice gas algorithm maintains Galilean invariance if we seek the sampling lattice gas algorithm to be competitive with the lattice Boltzmann algorithm.
To numerically test the Galilean invariance of our system, we initialized a set of SOD shock tube systems similarly to those in Fig. 8 for both the lattice Boltzmann and lattice gas algorithms with a lattice size of $L=1,000$ and a net velocity of zero, and a second set of identical systems with a non-zero constant velocity offset $u_{g}$. After running all simulations (including averaging 2,500 random seeds for each lattice gas system) for 400 iterations at $\omega=\omega_{LB}=1$, we adjusted the coordinate systems for the initial $u_{g}$ values to take into account the additional distance the simulation would have traveled on average due to the velocity by using,
@@
which takes into account our system’s periodic boundary conditions. $a]\bmod[b]$ is the modulo operator, returning the positive remainder when $b$ divides $a$. We then calculated the mean squared error in $\bar{N}$ and $u$,
@@
A comparison of the $\epsilon$ values for $\bar{N}$ and $u$ for both the lattice gas and lattice Boltzmann algorithms are shown in Fig. 14 (a) and 14 (b) for various values of $u_{g}$ ranging from 0.2-0.2 to $0.2$. The mean squared error for the lattice gas algorithm is on an order of magnitude comparable to that of the lattice Boltzmann algorithm for both $\bar{N}$ and $u$, meaning that the lattice gas and lattice Boltzmann algorithms are approximately equivalent in terms of Galilean invariance. There is a slight deviation observed for higher values of $u_{g}$, which implies that the lattice gas algorithm is slightly more Galilean invariant than the lattice Boltzmann algorithm, however we are as of yet unsure why this deviation occurs.
@@ V Computational Efficiency
V.1 Computational Efficiency Results
The main drawback of Blommel and Wagner’s algorithm was its lack of computational efficiency. This caused the method to remain non-competitive when compared to the contemporary fluctuating lattice Boltzmann. Thus, if we are to present our sampling lattice gas as a viable alternative to the fluctuating lattice Boltzmann for fluctuating systems, we must compare the computational efficiency of our algorithm to that of the lattice Boltzmann algorithm.
@@ V.1 Computational Efficiency Results
Figure 15: Computational efficiency of the lattice Boltzmann algorithm compared to various lattice gas algorithms for average system densities from $N=10$ to $N=1,000$. (a) compares the lattice Boltzmann algorithm to Blommel and Wagner’s algorithm, our sampling lattice gas algorithm, and a second run of our lattice gas algorithm with all randomly sampled numbers pre-generated for $\omega=\omega_{LB}=1$, leaving out Blommel and Wagner’s algorithm since it has no true way to reach $\omega=1$. (b) compares the lattice Boltzmann algorithm to our sampling lattice gas algorithm, Blommel and Wagner’s algorithm, and a second run of our lattice gas algorithm with all randomly sampled numbers pre-generated for $\omega=\omega_{LB}=0.5$ See text for further details.
Fig. 15 shows the results of the computational efficiency tests for four algorithms: the lattice Boltzmann discussed in Sec. III, Blommel and Wagner’s algorithm, a sampling lattice gas lookup table based algorithm discussed in the next section, and a sampling lattice gas with pre-generated random numbers. Specifically, the algorithms are compared based on the number of lattice updates they compute per second, given initial average densities ranging from $N=10$ to $N=1000$ average particles per lattice site. By timing how long it takes for the algorithm to run a particular number of iterations, then converting this time based on the lattice size, we can compare the number of lattice updates directly. For Fig. 15, these quantities are 10,000 timesteps per run and a lattice size of 100. Due to the high order of magnitude of the number lattice updates per second for the given algorithms, Fig. 15 has units of millions of lattice updates per second, or MLUPS.
It bears noting that the efficiency can vary wildly based on the specs of computer used for simulation. For consistency, all simulations were run on the same Dell XPS laptop with an Intel I5 CPU, and a clock speed of 1.6 GHz. Along with this, because of the way the lookup tables for the sampling lattice gas (which will be discussed further in the next section) work, the computational cost of generating them drops to zero as the simulation time increases, thus we run each lattice gas system for 10,000 timesteps prior to the 10,000 timesteps we time to minimize the cost of this generation. It is also worth noting that the entropic lattice Boltzmann algorithm we are using for comparison is not a fluctuating method, and thus is not truly comparable to the lattice gas system in terms of efficiency. A true fluctuating entropic lattice Boltzmann method would be less efficient than the algorithm used here. Analysis of the local equilibrium $\pi$ ensemble also suggests that for high density systems the distribution becomes approximately normal. Using this approximation to sample, as has been done by Chopard et al. [15, 16, sec. 5.7], may provide even more efficiency for our lattice gas than the explicit distribution sampling methods used here due to the relative efficiency of sampling from a normal distribution.
In Fig. 15 (a), the sampling lattice gas simulates an average of approximately 122.6 MLUPS for densities greater than or equal to 600 average particles per lattice site, and an average of approximately 175.8 MLUPS for densities less than 600 average particles per lattice site. The lattice Boltzmann simulates a relatively consistent average of approximately 162.8 MLUPS for all densities. Thus, for low densities the sampling lattice gas is more efficient than the lattice Boltzmann algorithm, with an average efficiency approximately 7.9% higher than the lattice Boltzmann. For high density systems, the lattice gas is still on the same order of magnitude as the lattice Boltzmann, with an efficiency loss of approximately 24.7%.
The $\omega=1$ case shown here does not include Blommel’s algorithm, due to the fact that to approximate a particular value of $\omega$ for Blommel’s algorithm requires that an associated number of collisions must be done to reach the appropriate number of collided particles on average. The number of collisions $C$ needed for a lattice site with $N$ particles to reach a particular value of $\omega$ is given by, [21]